OpenStudy (anonymous):

how do u get the answer . The sum of the sides of a triangle is represented by 20x. If one side is expressed by 8x and another side is expressed by 5x, what is the length of the third side? A) 13x B) 12x C) 7x D) 15x

OpenStudy (anonymous):

C

OpenStudy (anonymous):

5x +8x +7x=20x

OpenStudy (anonymous):

sum of 3 side = 20x sum of 2 side = 5x + 8x so what's sum of side left alone ??

OpenStudy (anonymous):

20x-5x+8x= ? 20x-13x=? ...7x

OpenStudy (anonymous):

Do you understand everything here?

OpenStudy (anonymous):

tha

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

sure anytime camo

OpenStudy (anonymous):

lol